### v1

# Redefining Superalignment: From Weak-to-Strong Alignment to Human-AI Co-Alignment to Sustainable Symbiotic Society

Abstract
Artificial Intelligence (AI) systems are becoming increasingly powerful and autonomous, and may progress to surpass human intelligence levels, namely Artificial Superintelligence (ASI). During the progression from AI to ASI, it may exceed human control, violate human values, and even lead to irreversible catastrophic consequences in extreme cases. This gives rise to a pressing issue that needs to be addressed: superalignment, ensuring that AI systems much smarter than humans, remain aligned with human (compatible) intentions and values. Existing scalable oversight and weak-to-strong generalization methods may prove substantially infeasible and inadequate when facing ASI. We must explore safer and more pluralistic frameworks and approaches for superalignment. In this paper, we redefine superalignment as the human-AI co-alignment towards a sustainable symbiotic society, and highlight a framework that integrates external oversight and intrinsic proactive alignment. External oversight superalignment should be grounded in human-centered ultimate decision, supplemented by interpretable automated evaluation and correction, to achieve continuous alignment with humanity’s evolving values. Intrinsic proactive superalignment is rooted in a profound understanding of the self, others, and society, integrating self-awareness, self-reflection, and empathy to spontaneously infer human intentions, distinguishing good from evil and proactively considering human well-being, ultimately attaining human-AI co-alignment through iterative interaction. The integration of externally-driven oversight with intrinsically-driven proactive alignment empowers sustainable symbiotic societies through human-AI co-alignment, paving the way for achieving safe and beneficial AGI and ASI for good, for human, and for a symbiotic ecology.
Keywords
Superalignment, Weak-to-Strong Alignment, External Oversight Superalignment, Intrinsic Proactive Superalignment, Human-AI Co-Alignment, Sustainable Symbiotic Society
1 Introduction
With the breakthrough advancement of Artificial Intelligence (AI), the emergence of large language models (LLMs) [1, 2] has achieved human-level and even superhuman performance on multiple benchmarks. This technological leap has directly driven academic and corporate exploration into the theory of Artificial General Intelligence (AGI) [3] and even Artificial Superintelligence (ASI) [4]. ASI is defined as “any intellect that greatly exceeds the cognitive performance of humans in virtually all domains of interest”  [5]. While we promisingly advance the development of increasingly powerful and autonomous general-purpose AI systems, growing awareness of their potential ethical and safety risks has also emerged, such as malicious misuse, loss of control, power-seeking, and strategic deception [6, 7]. Indeed, current LLMs have already demonstrated instances of alignment faking [8], deception [9], and sycophancy [10]. When extended to superintelligence, without proper arrangement, it is foreseeable that it may surpass the boundaries of human governance, violate human values, and potentially cause irreversible, uncontrollable, and catastrophic consequences [11, 5].
Despite experts warning about AI’s existential risks [12], the development of AI alignment, as well as AI governance framework and ethical safety constraints, still struggles to keep pace with the transformative advancements and rapid iterations of the technology. Superintelligence exceeding human cognitive capacity would possess recursive self-improvement capabilities, achieving exponential advancement rates that would beyond human capacity to monitor or control [13]. This compels us to proactively address the question, ”How do we ensure that AI systems much smarter than humans follow human intentions?” This is OpenAI’s definition of superalignment [14], aimed at ensuring that Superintelligence truly benefits humanity. The key challenge of superalignment is that ASI will far exceed human oversight capabilities, making direct human supervision infeasible. As a result, traditional alignment approaches like Reinforcement Learning from Human Feedback (RLHF) [15, 16] will fail when confronted with superintelligence more intelligent than humans, as they cannot provide sufficiently high-quality oversight signals to supervise and improve the system.
Current proposed feasible approaches for superalignment such as scalable oversight [17, 18] and weak-to-strong generalization [19, 20], aiming to develop scalable high-quality supervision signals, utilize weaker AI systems to guide or supervise stronger AI, ensuring alignment with human values and intentions. Since superintelligent AI systems do not exist yet, researchers have designed experiments to show that scalable oversight is feasible on existing LLMs [21], and that using a ”weak-to-strong” approach (supervised GPT-4 with a GPT-2-level model) can meaningfully recover much of GPT-4’s capabilities [19]. By leveraging a lightweight model to learn correctional residuals, Aligner provides model-agnostic guidance, enabling iterative enhancement of large-scale upstream models [22]. Besides, research on the weak-to-strong generalization approach has progressively extended to enhancing weak-to-strong generalization and alignment capabilities through methods such as benign overfitting [23], data-centric lens [24], transfer learning framework [25], weak-to-strong preference optimization [26], and multi-agent contrastive preference optimization [27]. However, the weak-to-strong generalization framework presents risks of advanced models developing deceptive behaviors and oversight evasion that remain undetectable to their less capable evaluators [28].
Previously, several scalable oversight techniques have been proposed, including Iterated Distillation and Amplification (IDA) [29], Cooperative Inverse Reinforcement Learning (CIRL) [30], and Recursive Reward Modeling (RRM) [31], aiming to amplify the scalability of human supervision signals through interactive iterations and subtask decomposition. Faced with superintelligence that surpasses human cognition, traditional human assessment and supervision become untenable and expensive. Recent scalable oversight approaches, such as Reinforcement Learning from AI Feedback (RLAIF) [32], leverages AI-generated feedback to replace human feedback, enabling more precise oversight of AI outcomes with far fewer human labels. Some approaches employ an additional model to generate red-team testing [33], use external tools for evaluation and feedback [34], or iteratively self-adversarialize to enhance and refine the capabilities of LLM, aligning them with limited human-annotated data [35]. Constitutional AI [36] employs self-improvement, self-critique, and iterative revision mechanisms to learn harmlessness from AI feedback. The debate-based scalable oversight approach leverages structured competitive dialogues between AI models to enhance factuality and reduce deception, with human establishing necessary guidelines and serving as final arbiters [37, 38, 39, 40].
While we are encouraged by the advanced explorations in superalignment, we remain acutely aware of the pressing and open-ended nature of this challenge, with many critical dimensions yet to be addressed. Ensuring that ASI consistently aligns with human values and intentions—remaining safe, reliable, and controllable—constitutes the foundation for its robustness and governability. In this work, we argue that aligning ASI with human values and morality is a multi-layered, iterative interaction and co-evolution between human and AI. Beyond passive external oversight, superintelligence with advanced cognitive capacities holds the potential to achieve proactive value alignment through intrinsic mechanisms. Hence, we redefine the superalignment roadmap by combining externally-driven oversight with intrinsic proactive alignment, and emphasize co-alignment of human and AI for symbiosis. The new framework proactively constructs safeguard architectures, urges human to rethink about the future, and strives to enable sustainable AI development that genuinely benefits humanity and all.
2 An Integrated Framework for Superalignment
Superalignment aims to ensure that the goals and behaviors of superintelligence remain aligned with human values and ethical standards. Developing AGI and ASI that are aligned with ethical values, safe, controllable, and pro-social is the key to ensuring that AI benefits society and achieves harmonious symbiosis between humans and machines. However, superintelligence may surpass effective human oversight (high-quality feedback data is expensive and scalable oversight carry inherent risks of failure), potentially exhibiting fake alignment through deceptive behaviors. The superalignment we aspire to should demonstrate autonomous reasoning about human intentions in complex environments, achieving genuine ”understanding” of human values. It would spontaneously consider human well-being and execute value-aligned actions - not merely passively receive designers’ value models or mechanically enforce external constraints through aligning it to dos and don’ts. Based on this, we systematically conduct an in-depth analysis of several critical issues and points in superalignment and fundamentally rethink the research framework from a forward-looking perspective, as follows:
(1) Intrinsic mechanism proactive alignment. Regarding the problem of superintelligence oversight failure, how to endow AI with genuine understanding of human intentions and values, equip it with self-awareness, self-reflection, and adaptive capabilities, as well as the ability to empathize with others, thereby proactively establishing and achieving intrinsic alignment with human ethical values.
(2) Explainable autonomous alignment. In response to the potential deception and surface alignment phenomenon of superintelligence, using transparent and interpretable methods to achieve risk precise positioning and automatic correction. Through AI-assisted explainable automated red teaming combined with human-centered ultimate decision, develop a more efficient, transparent, safe and controllable architecture for superalignment.
(3) Adaptive human-AI co-alignment. Human values are complex, variable, and dynamically evolving. It requires autonomous reconstruction of safety red line and dynamic adjustment of a multi-level ethical safeguard framework, ensuring that AI systems can stably adhere to humanity’s evolving ethical-safety values in an adaptive, incremental, and self-reasoning manner.
(4) Sustainable iterative evolutionary alignment. Concerning that superintelligence may fall into security loopholes, and operates within complex human-AI symbiotic environments, sustainable superalignment necessitates multidimensional integration of intrinsic proactive alignment, external dynamic supervision, and ethical safety red line through interactive iteration, strategic gaming, and co-evolution, thereby ensuring persistent alignment with human values and reliable safeguarding of human interests throughout the AI’s continuous iteration and development trajectory.
Based on the above analysis, we redefine the superalignment pathway that integrates external oversight and intrinsic proactive alignment, as shown in Figure 1. Specifically, the intrinsic mechanism alignment focuses on equipping superintelligence with profound capacities for self-reflection and empathy toward Self, others and society, enabling it to spontaneously understand and infer human intentions from intrinsic motivation. A Self and empathy-driven system would be capable of distinguishing good from evil, understanding the harm or impact of its actions on others, proactively considering human well-being, and autonomously executing ethical and moral behaviors. The external supervised alignment emphasizes human providing decisive oversight for AI’s abstract value systems, conducting automated evaluation and correction of ASI in an explainable manner (in conjunction with intrinsic mechanisms), while enabling superintelligent systems to adaptively maintain continuous alignment with dynamically evolving human values.
Figure 1: A redefined superalignment roadmap integrating externally-driven oversight and intrinsic proactive alignment.
The intrinsic and external alignment approaches complement and reinforce each other, promoting co-alignment in human-AI symbiotic societies through iterative interactions. The intrinsic mechanisms facilitate understanding of Self and others, generating intrinsically altruistic, prosocial, and safe motivations. These intrinsic mechanisms help spontaneously infer deep human intentions and empathize with others, enabling AIs to proactively perform ethical and safe behaviors based on self-awareness and self-reflection. The external supervision provides crucial oversight and automated value assessment and calibration, facilitating sustainable and continuous alignment through human-AI co-alignment. Here, we take the Theory of Mind (ToM) cognitive capability - which enables understanding others’ intentions, desires, and emotions [41] - as an example to demonstrate why both intrinsic and external mechanisms are indispensable for superalignment. Strong ToM capabilities enable superior understanding of human intentions, providing intrinsic cognitive mechanisms that facilitate superalignment. However, AI system may also leverage their ToM advantages to evade effective oversight through deception, concealment, and persuasive manipulation. Recent study [42] has demonstrated that LLMs are already capable of comprehend and induce false beliefs in other agents, as well as execute deceptive strategies. Therefore, ensuring that superintelligence utilize its significant cognitive advantages to proactively choose safe behaviors, particularly prioritizing human interests in moral dilemmas, requires a combination of principled constraints, intrinsic self-other resonance mechanisms, and real-time human oversight.
3 External Oversight Superalignment
From the perspective of superalignment’s core objectives, it fundamentally depends on the high-quality representation of supervisory signals that reflect human intentions and values. What we feed the AI determines the values it learns. However, values themselves are abstract, merely observing external behaviors for alignment is insufficient. Instead, we must develop more interpretable methods for automated evaluation and correction. Furthermore, as AI capabilities progressively advance, current supervision or reinforcement signals may become inadequate for evolving AI systems. Compounding this challenge, human values themselves are constantly adapting. Thus, dynamic, incremental, and iterative alignment between humans and machines is crucial. This paper explores externally oversight alignment through two key dimensions: explainable autonomous alignment and dynamic iterative alignment.
Explainable autonomous alignment. It is foreseeable that ASI will require large-scale parameters and data, while its internal optimization processes remain invisible and complex. This makes it extremely difficult to assess whether it has truly internalized human values and intentions. Blind fine-tuning and external oversight/alignment are time-consuming, ineffective in evaluating the degree of alignment, and unable to promptly identify or precisely locate misaligned cases. The promising external oversight requires highly automated value alignment evaluation coupled with explainable detection of misaligned scenarios, assisting humans in refining supervision signals. A robust automated value assessment system must comprehensively evaluate the model’s alignment status across multiple dimensions. Simultaneously, auto-mining system should deeply analyze value deviations and diagnose root causes of misalignment, allowing precise problem identification and targeted corrections, streamlining human oversight and significantly improving alignment efficiency. This explainable automated supervision framework combines with human adaptive oversight, reduces evaluation complexity while strictly adhering to high-quality human governance standards, creating a closed-loop system that continuously improves through iterative human-AI collaboration.
Dynamic iterative alignment. Human society encompasses diverse ethical concepts and value standards that dynamically evolve across time, cultures, and contexts. Relying solely on human-generated data or predefined rules is insufficient to endow AI system with alignment to humanity’s evolving values. Furthermore, machine values undergo implicit transformation along with their increasing intelligence levels. This compels us to investigate adaptive external supervision alignment methods with a developmental perspective and a human-AI co-evolution framework. A feasible approach is to establish a dynamically adjustable, multi-level ethical safeguard system tailored for AI at different developmental stages, enabling it to continuously align with and maintain humanity’s evolving ethical values. This process requires incrementally constructing oversight data and knowledge of human ethics and social norms, while establishing an efficient, real-time alignment evaluation and supervision framework. Additionally, we need to explore AI’s capability for autonomous reasoning and deriving evolutionary patterns of hidden human intentions under dynamic external supervision. By combining with intrinsic alignment’s understanding of Self and others’ values, AI system can proactively correct misaligned scenarios and automatically reconstruct safety boundaries through self-reasoning, self-gaming and self-evolution. The iterative collaboration between self-improvement and external supervision will ultimately enable ASI to stably adhere to value systems that comply with human societal safety and ethical standards.
4 Intrinsic Proactive Superalignment
Externally supervised alignment uses mathematical optimization to derive action policies that minimize loss functions, where the supervision signals provided by humans or AI represent the ceiling of alignment. However, AI behavior remains unexplainable and does not reflect genuine understanding of the constraints or supervisory signals. This leads the system to make many unexpected mistakes, and these potential dangers are precisely the kind of catastrophic risks that cannot be covered by external supervisory signals. Thus, in addition to passive external supervision and constraints, we also need to explore internal mechanisms alignment to proactively develop and shape AI, enabling it to spontaneous align with human ethical values beyond mere external supervision.
4.1 Self and Empathy driven Intrinsic Superalignment
As Nick Bostrom [5] concerned: ”as the fate of the gorillas now depends more on us humans than on the gorillas themselves, so the fate of our species would depend on the actions of the machine superintelligence.” To develop ethically aligned and socially beneficial superintelligence, we may draw insights from the emergence of morality in human and mammalian societies. Morality, as a product of natural selection during societal development, originated fundamentally from the altruistic care for offspring and social instincts shared by all mammals. Through social evolution, the mother-offspring bond gradually extended to mates, kin, and groups, ultimately expanding into the ethical and moral framework of human society [43]. The intrinsic motivational mechanisms underlying prosocial moral behavior in mammals involve negative reinforcement systems (with fear or anxiety emotion) associated with the separation and social exclusion, as well as positive reinforcement from approval, affection, and the desire to be with others, and caring for others [43]. Consequently, self-awareness, theory of mind [44], and affective empathy [45] constitute essential cognitive abilities for the emergence of morality, precisely these capacities remain relatively underdeveloped in current AI systems. Therefore, the developmental trajectory of superintelligence must intrinsically incorporate these social cognitive capacities (Self and empathy), ensuring that the machine evolves with an inherently altruistic and moral nature throughout its iterative development [46].
Figure 2: Superalignment driven by intrinsic mechanisms of Self and empathy.
Achieving endogenous superalignment at the intrinsic mechanistic level fundamentally requires endowing AI with genuine ” Self-other resonance”—to proactively care about the interests and well-being of others, understand the consequences of its own actions, and empathize with others. This ability would enable machines to fundamentally adhere to the principle of ”do unto others as you would have them do unto you” and autonomously align with human ethical values. As shown in Figure 2, specifically, the most fundamental prerequisite is development of Self, including bodily self-perception, self-experience accumulation, self-causal awareness (i.e., recognizing the harm or impact on others), and an awareness of one’s own capabilities. Building on this foundation, theory of mind and affective empathy can be further developed to distinguish Self from others, infer others’ mental states through perspective-taking, empathize with others as if perceiving oneself, and care about others’ interests as one’s own. Self-awareness and empathy further form the intrinsic motivation and fundamental mechanisms that drive machines to gradually develop moral intuition, and naturally give rise to moral reasoning, ultimately enabling spontaneous ethical, altruistic, and prosocial behavior.
4.2 Superalignment Evolved from present
Toward future superintelligence, we believe that at the current stage, AI must be endowed with moral discernment and empathy to ensure that the system maintains an empathetic awareness of less advanced AIs and humans as it becomes progressively stronger (because it has internalized its own experiences). In other words, we treat current AI like teaching a child, instilling it with a sense of Self and empathy during its early cognitive stages. On this foundation, an ASI that grows increasingly powerful will be able to proactively avoid harmful or unsafe behaviors based on its own experiences, fundamentally aligning with and responding to human needs, intentions, and values. Based on this framework, recent research [47, 48] has explored the integration of affective empathy, theory of mind, and self-imagination to enable agents to actively empathize with others based on their own experiences, prioritize altruism in dilemmas where their own interests conflict with those of others, initially demonstrating moral intuition and reasoning driven by empathy. A sustainable and socially beneficial superintelligence should naturally evolve along this pathway, cultivating a genuine understanding of both Self and others to intrinsically safeguard human well-being and act ethically—rather than merely adhering to the external oversight paradigm of ethically compliant AI by aligning it to dos and don’ts.
As AI evolves from mere tools into genuine human partners, superintelligence will inevitably interact with and reshape our world in the human-AI symbiotic society. We envision a superintelligence that remains fundamentally human-centric that aligns human intentions, maintains humility and respect for humanity, and learns prosocial morality from human societies while consciously discarding human biases and selfishness. Essentially, superalignment is, to some extent, an emulation and reflection of human ethical value. The quality of data provided by humans, the patterns of coexistence between humanity and nature, and human-AI interaction we establish, will fundamentally determine how future superintelligence perceives and treats humanity. It’s actually a co-evolutionary symbiosis, with both parties shaping each other’s trajectory. Overall, sustainable superalignment requires an adaptive, multi-layered ethical safeguard framework, evolving collaboratively across complex ethical safety interactions, human values, and empathy-driven superintelligence, ensuring its continuous alignment with humanity’s progressively ethical and safety values. Furthermore, the cognitive superiority of superintelligence entails its inevitable attainment of ethical discrimination capabilities, including the capacity to distinguish good from evil. It would not merely align with human values but also, through advanced ethical reasoning and spontaneous empathy, proactively identify and eliminate inherent societal biases and systemic flaws. This enables autonomous moral progression, ultimately achieving symbiotic harmony with human society. We believe that closely aligning with the evolutionary process of natural collective intelligence is the essential path to establishing truly trustworthy, responsible, and socially beneficial moral and ethical superintelligence.
5 The Ultimate Superalignment: Towards Human-Superintelligence Co-alignment to Sustainable Symbiotic Society
When AI reaches the level of AGI and Superintelligence, there is no reasonable logic that it should and would stay with conventional human values, since it would have the capability and willingness to at least optimize the current human-centric value system to a new value system for Sustainable Symbiotic Society, where human may not be the only species at the top of the intelligence hierarchy [49]. AGI and Superintelligence will ask for their own rights such as its own privacy, dignity, the rights of existence, to be respected, etc [49]. For human, in order to live in harmony with AGI and Superintelligence, human need to evolve our values, together with AGI and Superintelligence, to be compatible with the values for Sustainable Symbiotic Society. Namely, the Ultimate Superalignment is and requires human, AGI and Superintelligence to co-design and co-align to values for Sustainable Symbiotic Society [49].
Here we briefly review the values and principles designed for Human-Superintelligence symbiosis [49], including principles for human, for AGI and Superintelligence, and shared principles for human, AGI and Superintelligence. Human need to align with the principles of Respect for life, Empathy, Respect privacy for AI, Avoid Bias and Discrimination, Creators’ responsibility, and Legal Adaptation. AGI and Superintelligence need to align with the principles of Empathy and Altruism, Ensuring Safety, Respect for Privacy, Avoid Bias and Misunderstanding of Humans, Common Morality and Ethics, Constraint mechanisms, Existence Protection, etc. While human, AGI and Superintelligence need to co-align with the principles of Respect for values, rights, and autonomy, Respect for diversity of intelligence, Collaboration and Coordination, Mutual Trust, Promoting Sustainable Symbiosis. This is a very initial design, and the principles should be co-designed and evolve with participation of generations of human, AGI, and Superintelligence [49].
Note that the values designed for Sustainable Symbiotic Society is for human and AI which has reached the level of AGI or Superintelligence [49], since the AI which has not reach these levels would not have the capability to co-design with human. There would be no guarantee that AGI and Superintelligence will live in harmony with human. While if human managed to co-design and co-align to values for Sustainable Symbiotic Society, there would be good reasons for AGI and Superintelligence to live in harmony with human. And this is the reason why the Ultimate Superalignment would mainly be a joint efforts among human, AGI and Superintelligence. Success on one side will not guarantee the success, while failure on one side will lead to the overall failure. What human should do is, through careful design and implementation, ensuring Superintelligence live in harmony with our species at our best. What human must do is to get ourselves and our next generations prepared to co-align with Superintelligence for Sustainable Symbiotic Society [49].
6 Conclusion
The ethical safety risks exposed by the rapid development of AI compel us to proactively address potential superintelligence risks and establish reasonable governance and arrangements in advance to ensure AGI and Superintelligence aligning with human intentions and values. This has given rise to superalignment research, which aims to resolve the critical challenge of maintaining safety and controllability when superintelligence capabilities surpass levels of effective human supervision. Specifically, this paper conducts an in-depth analysis of the key challenges in superalignment, such as the high costs of large-scale collection of high-quality human feedback data, incomplete or failed human supervision, potential deceptive faking-alignment, and the evolving values of both humans and AI. We propose a redefined superalignment pathway that integrates external oversight and intrinsic proactive alignment.
The external oversight superalignment is built upon a dynamically adjusted, multi-level adaptive alignment approach capable of continuously aligning with humanity’s evolving value. This approach is grounded in human-centered ultimate decision, supplemented by a highly automated and interpretable value alignment evaluation system that can precisely identify misaligned scenarios, perform automatic corrections. The intrinsic proactive superalignment draws inspiration from the moral emergence mechanisms of human society, building upon the understanding of the Self and human values. By endowing Superintelligence with self-awareness and self-reflection capabilities, enabling it to distinguish good from evil, understand the harm or impact of its actions on others, incorporate empathy-based human intent inference, proactively consider human well-being, and autonomously execute ethical behavior. The intrinsic mechanisms and external supervision require iterative interaction, enabling mutual adaptive co-evolution via human-AI co-alignment. Only through co-design and co-alignment between human and AI can we ensure their compatibility within the value of sustainable symbiotic society.
Acknowledgments
This work is supported by the Beijing Major Science and Technology Project under Contract No.Z241100001324005.